OpenStudy (hiyathere_20):

Write the equation of the quadratic function with roots -1 and -7 and a vertex at (-4, 7)

OpenStudy (mathmale):

Use the vertex form of the equation of a parabola (quadratic function) as a starting point:\[y-k=a(x-h)^2\] where (h,k) is the vertex.

OpenStudy (mathmale):

Alternatively, you could use y=ax^2 + bx + c. In this case, the x-coordinate of the vertex is \[x _{vertex}=\frac{ -b }{ 2a }\]